OpenStudy (anonymous):

SOME ONE PLEASE HELP The steps to derive the quadratic formula are shown below: Step 1: ax2 + bx + c = 0 Step 2: x2 + b over ax + c over a = 0 Step 3: x2 + b over a x = - c over a Step 4: x2 +b over a x + square of quantity of b over 2 times a = -c over a + square of quantity of b over 2 times a Step 5: x plus b over 2 times a, all squared = -c over a + square of quantity of b over 2 times a Step 6: x + b over 2 times a = square root of all of negative c over a plus the square of the quantity b over 2 times a Step 7: x = - b over 2 times a ± square root of all of negative c over a plus

OpenStudy (anonymous):

Which step is wrong

OpenStudy (anonymous):

I don't see anything wrong. You do need to finish.\[(x+\frac{ b }{ 2a })^{2}=-\frac{ c }{ a} +(\frac{ b }{ 2a })^{2}\]\[x+\frac{ b }{ 2a }=\pm \sqrt{-\frac{ c }{ a}+(\frac{ b }{ 2a })^{2}}\]\[x=-\frac{ b }{ 2a }\pm \sqrt{-\frac{ c }{ a }+(\frac{ b }{ 2a })^{2}}\]Finishing \[x=-\frac{ b }{ 2a }\pm \sqrt{\frac{ b ^{2} }{ 4a ^{2} }-\frac{ 4ac }{ 4a ^{2} }}\]\[x=-\frac{ b }{ 2a }\frac{ \pm \sqrt{b ^{2}-4ac} }{ \pm \sqrt{4a ^{2}} }\]\[x=-\frac{ b }{ 2a }\pm \frac{ \sqrt{b ^{2}-4ac} }{ 2a }\]